# Measuring Computational Universality of Fully Homomorphic Encryption

### v1

@@ 2.2. Universality Required Privacy-sensitive Applications
Over the past decade, FHE has evolved from a purely theoretical construct into a practical solution for today’s data-centric challenges (Gouert et al., 2023). In finance, Intesa Sanpaolo’s collaboration with IBM leverages FHE secure asset transactions (IBM, 2024), allowing for both confidential database querying and the execution of complex machine learning algorithms on encrypted data. Similarly, in the medical domain, Roche’s partnership with ETH investigates the use of FHE for secure graph analyses of genomic information (Lou, 2024; Zama, 2022d). As industries increasingly adopt advanced analytics that require both linear and non-linear operations, FHE’s universal capacity to handle complex computations on encrypted data makes it an indispensable tool for ensuring data confidentiality and regulatory compliance.
@@
In Section 3, we analyze the universality of various FHE methods from a theoretical perspective. Section 5 presents experimental evaluations assessing their performance across diverse workloads. These workloads include sequences of operations such as non-linear applied after linear operations, mirroring the activation functions in neural networks; as well as linear following non-linear operations, which form the basic unit of fundamental $\text{Max}/\text{Min}(A,B)$ operations, expressed as $\text{Compare}(A,B)\times A+(1-\text{Compare}(A,B))\times B$. Finally, in Section 6, we evaluate privacy-sensitive applications built on these workloads, demonstrating the practical universality of FHE methods. The applications under study include graph algorithms, decision trees, database aggregation, neural-network inference, and sorting, each representing a critical function in real-world research and commercial settings. Benchmarking different FHE methods on these tasks provides insight into their real-world performance in universal scenarios.
3. FHE Universality Taxonomy and Complexity Analysis
@@
Table 1. Universality comparison of FHE methods under $b$-bit precision and $2^{\lambda}$ security level. N is the ring dimension. For a polynomial approximation of CKKS, the $n$ denotes the degree of evaluation polynomial.
Methods Non-linear Operation (i.e., Comparison) Linear Operation (i.e., Multiplication) Universal
@@
Word-wise BGV/BFV (Brakerski et al., 2014; Brakerski, 2012) - - - - - $1$ $O(\lambda^{2})$
Word-wise CKKS (Cheon et al., 2017) - - - - - $1$ $O(N\log N)$
Poly. Approx. (CKKS) (Cheon et al., 2020; Lee et al., 2021) $O(\log_{2}n)$ $O(\sqrt{n})$ $1$ $O(N\log N)$
Poly. Interp. (BGV/BFV) (Narumanchi et al., 2017) $O(b)$ $O(\sqrt{2^{b}})$ $1$ $O(\lambda^{2})$
Decomp. (BGV/BFV) (Iliashenko and Zucca, 2021) $O(\log_{2}b)$ $O(b\log_{2}b)$
XCMP (BGV/BFV) (Lu et al., 2018) $O(1)$ $O(1)$
General Functional Bootstrapping (Alexandru et al., 2024) $O(b)$ $O(\sqrt{2^{b}}+b)$ $1$ $O(N\log N)$
Bit-wise TFHE (Chillotti et al., 2020) $O(b)$ $O(b)$ $O(b^{2})$ $O(b^{2}p)\textit{CMUX}+O(bp)\textit{KS}$
Scheme Switching (Lu et al., 2021; Boura et al., 2020) - $\Omega(2^{b})$ $1$ $O(\lambda^{2})$
Encoding Switching (Zhang et al., 2023) $O(\log_{2}d+d\log_{2}p)$ $O(d^{2}\sqrt{p})$ $O(1)$ $O(\lambda^{2})$
@@
Universality is an important and general feature of current real-world privacy-sensitive applications, even though there is little research investigating the universality of various FHE methods. Specifically, a universal FHE method must satisfy two conditions:
@@
Mixed Linear and Non-Linear Operations: The universal FHE method must support both linear and non-linear operations on ciphertexts and allow them to be composed in any sequence. This enables complex computations that alternate between operation types, such as machine learning algorithms that use linear operations (e.g., matrix multiplications) followed by non-linear activation functions (e.g., ReLU). By seamlessly mixing linear and non-linear operations, universal FHE can accommodate various applications.
@@
Non-Interactive Computation: A universal FHE method should support mixed operations in a non-interactive manner—allowing computations on encrypted data without further communication between the data owner and the computing party. This non-interactivity is essential for scalability and efficiency, especially in distributed or cloud settings where reducing communication overhead is critical.
In this section, we evaluate the universality of prevalent FHE methods. We first classify them into two categories: universal FHE methods supporting both linear and non-linear operations, and non-universal FHE methods, which do not. We then provide a detailed analysis of each category in Sections 3.1 and 3.2, respectively, and summarize their universality properties and complexities in Table 1.
3.1. Analysis of universal FHE methods
@@
This section analyzes three universal FHE methods that support both linear and non-linear operations in a non-interactive manner. The first is bit-wise FHE (Chillotti et al., 2020; Ducas and Micciancio, 2015), which inherently supports both operation types by constructing arithmetic and Boolean circuits. The second is scheme switching (Lu et al., 2021; Boura et al., 2020), which combines bit-wise and word-wise FHE to leverage the strengths of both. The third is encoding switching (i.e., HEBridge (Zhang et al., 2023)), which integrates linear and non-linear operations within word-wise BGV/BFV methods by bridging distinct encoding spaces using proposed homomorphic reduction and lifting functions.
@@
Bit-wise FHE. TFHE (Chillotti et al., 2020) is a bit-wise FHE method that supports gate-level operations on encrypted data, enabling fast non-linear operations due to its efficient bit-wise processing. For a ciphertext with $b$-bit precision, TFHE can perform non-linear functions with a depth of $O(b)$ and complexity of $O(\sqrt{2^{b}})$. However, TFHE’s linear operations are comparatively slower due to its bit-wise design. Multiplication on ciphertexts requires a depth of $O(b^{2})$, along with $O(b^{2}p)$ CMUX operations and $O(bp)$ key-switching. Additionally, TFHE lacks support for batch processing, i.e., SIMD, making linear computations less efficient than in word-wise FHEs, which can parallelize linear operations.
@@
Scheme Switching. Scheme switching (Boura et al., 2020; Lu et al., 2021) between bit-wise and word-wise ciphertexts enables universal FHE operations by leveraging the strengths of both schemes: linear functions are evaluated in word-wise FHE, and non-linear functions in bit-wise FHE. However, the complexity of scheme switching grows at least exponentially with input bit-width (Ren et al., 2022; Bian et al., 2023), i.e., $\Omega(2^{b})$. For example, evaluating 6-bit inputs takes 43.8 seconds, while 8-bit inputs require 162.7 seconds. Consequently, scheme switching becomes impractical for larger inputs.
@@
Encoding Switching. Prior studies (Iliashenko and Zucca, 2021; Tan et al., 2020) have shown that word-wise BGV/BFV support efficient and precise non-linear operations via polynomial interpolation. However, these methods cannot be seamlessly incorporated with linear operations because the plaintext space for linear operations in FV over $\mathbb{Z}_{p^{d}}$ is different that used for non-linear operations in base-encoded FV (beFV) over $\mathbb{Z}_{p}$. To enable continuous evaluation of both operation types, Zhang et al. (Zhang et al., 2023) introduced conversion techniques that employ a reduction function, which homomorphically converts the plaintext space from FV to beFV, allowing non-linear operations to follow linear operations, and a lifting function that restores the result from beFV back to the original FV space for subsequent linear operations. In terms of complexity, given that the FV plaintext modulus is $p^{d}$, the reduction function requires $O(d^{2}\sqrt{p})$ multiplications and $O(d\log_{2}p)$ multiplicative depth, while the lifting function requires $O(\sqrt{dp})$ multiplications and a depth of $O(\log_{2}d+\log_{2}p)$.
3.2. Analysis of un-Universal FHE methods
@@
Word-wise FHEs, including BGV, BFV, and CKKS, na"ıvely support efficient linear operations with SIMD. However, computing non-linear functions in SIMD-enabled word-wise FHE is non-trivial. As a result, many efforts have been made to enable non-linear functions within word-wise FHE to leverage SIMD. For the CKKS scheme, the most common approach is polynomial approximation (Cheon et al., 2020, 2019; Lee et al., 2021). On the other hand, for integer-based BGV and BFV, Narumanchi et al. (Narumanchi et al., 2017) proposed evaluating non-linear functions via polynomial interpolation over the base field.
CKKS with Polynomial Approximation. CKKS is a word-wise FHE method capable of encrypting real numbers. Since it supports approximate computation on floating-point numbers, the most common method for performing non-linear functions is polynomial approximation (Lee et al., 2023b, 2022a). Specifically, these methods use compositions of minimax approximation polynomials to approximate the sign function over a small interval, typically $[-1,-\epsilon]\cup[\epsilon,1]$, with errors. Given an $n$-degree approximation polynomial, the complexity is $O(\sqrt{n})$ and the multiplicative depth is $\log_{2}n$. The larger the interval and the smaller the error, the higher the polynomial degree required—thus increasing the number of multiplications and multiplicative depth. Most importantly, approximation can never be accurate around 00. These errors limit the applicability in domains where accuracy is critical, such as genomics (Kim and Lauter, 2015; Raisaro et al., 2018; Zhang et al., 2015) and finance (Armknecht et al., 2015; Han et al., 2019).
BGB/BFV with Polynomial Interpolation. In contrast to CKKS, computations in BFV/BGV are exact. Accordingly, non-linear functions can be implemented in BFV/BGV without approximation error. Narumanchi et al. (Narumanchi et al., 2017) proposed computing non-linear functions via polynomial interpolation.
Take the non-linear comparison function as an example, since the comparison is a fundamental operation for many logic functions and is key to evaluating various non-linear operations (Zhang et al., 2023). The core idea is to convert the comparison between two encrypted integers, $a$ and $b$, into a comparison between their difference, $z=b-a$, and zero. This is achieved by constructing a polynomial $P(x)$ with the property:
$P(x)$ are expressed using Lagrange interpolation:
(3) $\text{Enc}(z)=P(\text{Enc}(a-b)).$
The polynomial interpolation has notable limitations. Specifically, the standard BGV and BFV support efficient linear operations over $\mathbb{Z}_{p^{r}}$ (Chen and Han, 2018; Halevi and Shoup, 2021). However, polynomial interpolation in $\mathbb{Z}_{p^{r}}$ is impractical, as it leads to polynomials of degree $O(p^{r})$, which grow exponentially with the input bit-width. Evaluating such high-degree polynomials is prohibitive in practice (Iliashenko and Zucca, 2021; Tan et al., 2020). Therefore, non-linear operations are typically performed in the base field $\mathbb{Z}_{p}$. The degree of interpolation polynomials equals $p$, requiring $O(\sqrt{p})$ non-scalar multiplications and $O(\log_{2}p)$ multiplicative depth using the Paterson-Stockmeyer algorithm. To enable efficient evaluation, the plaintext space is typically confined to a small prime $p$, making it unsuitable for large inputs. The value of $p$ ranges from $2\sim 7$ in (Tan et al., 2020), is at most $173$ in (Iliashenko and Zucca, 2021) and at most $257$ in (Morimura et al., 2023). In conclusion, this limitation makes it impractical for real-world universal applications.
Polynomial Interpolation with Special Encoding. Aiming to enhance the capability of polynomial interpolation to support non-linear operations on large inputs, several special encoding-based methods have been proposed (Iliashenko and Zucca, 2021; Tan et al., 2020), which encode a large integer as a vector of base-$p$ digits: $a=(a_{0},a_{1},\ldots,a_{r-1})$, where each $a_{i}\in\mathbb{Z}_{p}$. Consequently, a non-linear operation, e.g., a comparison between two integers $a$ and $b$, is transformed into a lexicographical comparison of their respective digit vectors. Formally, the process begins with the decomposition of the input integers $a$ and $b$ into their digit representations:
(5) $\text{Enc}(z_{i})=P(\text{Enc}(a_{i}-b_{i})).$
(7) $\log_{2}\log_{p}2^{b}+\log_{2}(p-1)+4.$
In contrast, directly interpolation on $\mathbb{Z}_{p^{r}}$ has a multiplicative depth of $O(\log_{2}p^{r})$. This reduced depth makes the method more practical for larger fields. By addressing the challenge of deep circuit depth in polynomial interpolation, (Iliashenko and Zucca, 2021) enables faster homomorphic non-linear operations for practical applications in BGV/BFV. Despite its significant efficiency improvement, the special encoding-based ciphertexts do not support linear operations since the number is expressed in the vector form, making it not a universal scheme.
Similarly, other special encoding-based approaches have also been proposed to enhance the efficiency of non-linear operations in word-wise FHE. Unfortunately, these methods sacrifice universality, as the specialized ciphertext formats do not support linear operations. Below, we present a representative example.
Exponential Encoding (XCMP). The exponential encoding method for FHE private comparison, named XCMP, was first proposed by Lu et al. (Lu et al., 2018). The core idea is based on the fact that the message space in HE is a polynomial ring, i.e., $\mathbb{Z}_{p}[x]/\langle x^{n}+1\rangle$, and that ciphertext multiplication corresponds to polynomial multiplication. This enables the design of a special encoding scheme that encodes the values to be compared into the polynomial’s degree coefficients.
The comparison process of integers $a$ and $b$ begins with encoding $a$ and $-b$ as polynomials $X^{a}$ and $X^{-b}$, respectively. These polynomials are encrypted into ciphertexts $\text{Enc}(X^{a})$ and $\text{Enc}(X^{-b})$, which are used to perform:
  • If $a\leq b$, $X^{b-a}$ from $T(X)$ aligns with $X^{a-b}$, yielding $C(X)[0]=1$.
  • If $a>b$, the $(n-(a-b))$-th term of $T(X)$ aligns with $X^{a-b}$, resulting in $C(X)[0]=-1$ due to the wrap-around: $X^{n-(a-b)}\times X^{a-b}\equiv-1\mod(X^{n}+1)$.
In other words, $C(X)[0]=1$ if $a\leq b$, and $C(X)[0]=-1$ if $a>b$.
XCMP has a multiplicative depth and complexity of 1 due to its use of a single ciphertext multiplication. However, it is limited by a small input field, requiring both $a$ and $b$ to be smaller than the polynomial degree $n$. This constraint reduces efficiency for large inputs, as a larger $n$ is needed (Lu et al., 2018). Although there have been discussions about extending the domain to $\mathbb{F}_{n^{2}}$ by decomposing numbers in base $n$ and processing digits separately:
where the equality check result $\mathbb{I}[a_{1}=b_{1}]$ must be converted to the XCMP format for multiplication with $\mathbb{I}[a_{0}\leq b_{0}]$. This conversion involves Fermat’s little theorem, resulting in a large multiplicative depth. As the result, XCMP performs well for small domains (less than 16 bits). For larger domains, performance significantly decreases (Lu et al., 2018).
General Functional Bootstrapping. Functional bootstrapping (Alexandru et al., 2024; Lee et al., 2024; Liu and Wang, 2023) extends traditional bootstrapping techniques by not only refreshing ciphertexts but also by enabling the homomorphic evaluation of arbitrary functions on encrypted data. The key idea is to express a target function as a look-up table (LUT) and then approximate this LUT via a constructed interpolation polynomial. Lee et al. (Lee et al., 2024) proposed functional bootstrapping for BFV, building upon regular BFV-style bootstrapping (Geelen and Vercauteren, 2023). However, BFV-style bootstrapping efficiently supports only a small number of slots. To improve efficiency, Alexandru et al. (Alexandru et al., 2024) employed CKKS-style bootstrapping to increase slot utilization. Specifically, the target function is first encoded as a LUT over a finite base-$p$ field $\mathbb{Z}_{p}$, and then approximated using a trigonometric Hermite interpolation polynomial $R(x)$. For instance, in the first-order case, the interpolation is constructed so that:
This polynomial is then evaluated homomorphically using an efficient technique such as the Paterson–Stockmeyer algorithm. The complexity of evaluating a polynomial of degree $d$ using this method is roughly proportional to $\sqrt{2d}+\log d$. For a first-order interpolation, where $d=p-1$, the complexity increases approximately as $\sqrt{p}$. Moreover, if further noise reduction is required, higher-order interpolations can be used (e.g., second-order or third-order), which increase the polynomial degree (to about $\tfrac{3p}{2}$ or $12p-$ 1, respectively) and thus the overall cost. Moreover, extending the input space from $\mathbb{Z}_{p}$ to $\mathbb{Z}_{p^{r}}$ requires representing each integer as a vector of base-$p$ digits, similar to special encoding-based polynomial interpolation in BFV/BGV (Zhang et al., 2023; Iliashenko and Zucca, 2021), which precludes the continuous evaluation of linear and non-linear functions on large inputs. Experiments in (Alexandru et al., 2024) demonstrate that general functional bootstrapping remains practical only for LUT evaluations up to 12-bit inputs.
4. FHE Libraries Supporting Universality

### v2

@@ 2.3. Generality Required Privacy-sensitive Applications
Over the past decade, FHE has evolved from a purely theoretical construct into a practical solution for today’s data-centric challenges (gouert2023sok). In finance, Intesa Sanpaolo’s collaboration with IBM leverages FHE secure asset transactions (finance_exp), allowing for both confidential database querying and the execution of complex machine learning algorithms on encrypted data. Similarly, in the medical domain, Roche’s partnership with ETH investigates the use of FHE for secure graph analyses of genomic information (lou2024homomorphic; hipaa). As industries increasingly adopt advanced analytics that require both linear and non-linear operations, FHE’s generality to handle complex computations on encrypted data makes it an indispensable tool for ensuring data confidentiality and regulatory compliance.
@@
In Section 3, we analyze the computational generality of various FHE methods from a theoretical perspective. Section 6 presents experimental evaluations assessing their performance across diverse workloads. These workloads include sequences of operations such as non-linear applied after linear operations, mirroring the activation functions in neural networks; as well as linear following non-linear operations, which form the basic unit of basic $\text{Max}/\text{Min}(A,B)$ operations, expressed as $\text{Compare}(A,B)\times A+(1-\text{Compare}(A,B))\times B$. Finally, in Section 7, we evaluate privacy-sensitive AI-related applications built on these workloads, demonstrating the practical generality of FHE methods. The applications under study include graph algorithms, decision trees, database retrieval, neural-network inference, and sorting, each representing a critical function in real-world research and commercial settings.
3. FHE Generality Taxonomy and Complexity Analysis
@@
Table 1. Functionality and cost comparison of representative FHE methods categorized by functional capability and theoretical complexity under $b$-bit precision and $2^{\lambda}$ security level. The last three lines represent the methods achieve full computational generality across linear and non-linear operations. NN is the ring dimension; for polynomial approximation of CKKS, $n$ denotes the polynomial degree.
Methods Non-linear Operation (i.e., Comparison) Linear Operation (i.e., Multiplication) Generality
@@
Word-wise BGV/BFV (brakerski2014leveled; brakerski2012fully) - - - - - 11 $O(\lambda^{2})$
Word-wise CKKS (cheon2017homomorphic) - - - - - 11 $O(N\log N)$
Poly. Approx. (CKKS) (cheon2020efficient; lee2021minimax) $O(\log_{2}n)$ $O(\sqrt{n})$ 11 $O(N\log N)$
Poly. Interp. (BGV/BFV) (narumanchi2017performance) $O(b)$ $O(\sqrt{2^{b}})$ 11 $O(\lambda^{2})$
Decomp. (BGV/BFV) (iliashenko2021faster) $O(\log_{2}b)$ $O(b\log_{2}b)$ - - - -
XCMP (BGV/BFV) (lu2018non) $O(1)$ $O(1)$ - - - -
General Functional Bootstrapping (alexandru2024general) $O(b)$ $O(\sqrt{2^{b}}+b)$ 11 $O(N\log N)$
Bit-wise TFHE (chillotti2020tfhe) $O(b)$ $O(b)$ $O(b)$ $O(b^{2}p)\textit{CMUX}+O(bp)\textit{KS}$
Scheme Switching (lu2021pegasus; boura2020chimera) - $\Omega(2^{b})$ 11 $O(\lambda^{2})$
Encoding Switching (zhang2023hebridge) $O(\log_{2}d+d\log_{2}p)$ $O(d^{2}\sqrt{p})$ $O(1)$ $O(\lambda^{2})$
@@
General computation is a crucial requirement of real-world privacy-sensitive applications, yet little research has investigated how existing FHE methods support this functionality. Specifically, an AI-friendly FHE method must satisfy two conditions:
@@
Mixed Linear and Non-Linear Operations: An FHE method with computational generality must support both linear and non-linear operations on ciphertexts and allow them to be composed in arbitrary sequences. This capability enables complex computations that alternate between operation types, such as machine learning algorithms that apply linear operations (e.g., matrix multiplications) followed by non-linear activation functions (e.g., ReLU). By seamlessly mixing linear and non-linear operations, such FHE methods can accommodate a wide range of applications.
@@
Non-Interactive Computation: An FHE method with computational generality should support these mixed operations in a non-interactive manner, allowing computations on encrypted data without further communication between the data owner and the computing party to decrypt and re-encrypt. Non-interactivity is essential for scalability and efficiency, especially in settings where minimizing communication overhead is critical.
Among existing FHE methods, only a few can simultaneously satisfy both of above conditions that can support both linear and non-linear computation in a non-interactive manner, what we term full computational generality. This section primarily focuses on these three computationally general FHE methods: bit-wise TFHE, Scheme Switching, and Encoding Switching. We analyze their theoretical complexities and evaluate their suitability for general AI computation in Section 3.1.
Other FHE methods, such as native word-wise BGV/BFV/CKKS, polynomial approximation, polynomial interpolation, and functional bootstrapping, provide only partial or approximate forms of computational generality. We discuss these method in Section 3.2 for conceptual completeness and to contextualize our focus. Table 1 summarizes the generality functionality and computational complexities of representative FHE methods, highlighting the distinction between general and un-general general approaches.
3.1. Analysis of General FHE methods
@@
This section analyzes three FHE methods that achieve computational generality, supporting both linear and non-linear operations in a non-interactive manner. The first is bit-wise FHE(chillotti2020tfhe; ducas2015fhew), which inherently enables both operation types through arithmetic and boolean circuits. The second is scheme switching(lu2021pegasus; boura2020chimera), which combines bit-wise and word-wise FHE to leverage the strengths of each. The third is encoding switching (i.e., HEBridge (zhang2023hebridge)), which integrates linear and non-linear operations within word-wise BGV/BFV by bridging distinct encoding spaces through homomorphic reduction and lifting functions.
@@
Bit-wise FHE. TFHE (chillotti2020tfhe) is a bit-wise FHE method that supports gate-level operations on encrypted data, enabling fast non-linear operations due to its efficient bit-wise processing. For a ciphertext with $b$-bit precision, TFHE can perform non-linear functions with a depth of $O(b)$ and complexity of $O(\sqrt{2^{b}})$. However, TFHE’s linear operations are comparatively slower due to its bit-wise design. Multiplication on ciphertexts requires a depth of $O(b^{2})$, along with $O(b^{2}p)$ CMUX operations and $O(bp)$ key-switching. Additionally, TFHE lacks support for batch processing, i.e., SIMD, making linear computations less efficient than in word-wise FHEs, which can parallelize linear operations.
@@
Scheme Switching. Scheme switching (boura2020chimera; lu2021pegasus) between bit-wise and word-wise ciphertexts enables general computation by leveraging the strengths of both schemes: linear functions are evaluated in word-wise FHE, and non-linear functions in bit-wise FHE. However, the complexity of scheme switching grows at least exponentially with input bit-width (ren2022heda; bian2023he3db), i.e., $\Omega(2^{b})$. For example, evaluating 6-bit inputs takes 43.8 seconds, while 8-bit inputs require 162.7 seconds. Consequently, scheme switching becomes impractical for larger inputs.
@@
Encoding Switching. Prior studies (iliashenko2021faster; tan2020efficient) have shown that word-wise BGV/BFV support efficient and precise non-linear operations via polynomial interpolation. However, these methods cannot be seamlessly incorporated with linear operations because the plaintext space for linear operations in FV over $\mathbb{Z}_{p^{d}}$ is different that used for non-linear operations in base-encoded FV (beFV) over $\mathbb{F}_{p}$. To enable continuous evaluation of both operation types, Zhang et al. (zhang2023hebridge) introduced conversion techniques that employ a reduction function, which homomorphically converts the plaintext space from FV to beFV, allowing non-linear operations to follow linear operations, and a lifting function that restores the result from beFV back to the original FV space for subsequent linear operations. In terms of complexity, given that the FV plaintext modulus is $p^{d}$, the reduction function requires $O(d^{2}\sqrt{p})$ multiplications and $O(d\log_{2}p)$ multiplicative depth, while the lifting function requires $O(\sqrt{dp})$ multiplications and a depth of $O(\log_{2}d+\log_{2}p)$.
3.2. Analysis of un-General FHE methods
While the three methods analyzed in Section 3.1 provide full computational generality, several other FHE approaches offer partial or approximate support for mixed linear and non-linear operations. We provide a brief overview of these methods here for conceptual completeness, emphasizing why they fall short of full generality.
@@ 3.2. Analysis of un-General FHE methods
Word-wise FHE Schemes. Word-wise FHEs, including BGV, BFV, and CKKS, naïvely support efficient linear operations with SIMD. However, computing non-linear functions in SIMD-enabled word-wise FHE is non-trivial. As a result, many efforts have been made to enable non-linear functions within word-wise FHE to leverage SIMD. For the CKKS scheme, the most common approach is polynomial approximation (cheon2020efficient; cheon2019numerical; lee2021minimax). On the other hand, for integer-based BGV and BFV, Narumanchi et al. (narumanchi2017performance) proposed evaluating non-linear functions via polynomial interpolation over the base field.
CKKS with Polynomial Approximation. CKKS can approximate non-linear functions (e.g., sign, ReLU, sigmoid) using compositions of minimax approximation polynomials over small intervals (lee2023precise; lee2022privacy) such as [1,1][-1,1]. Given an $n$-degree approximation polynomial, the complexity is $O(\sqrt{n})$ and the multiplicative depth is $\log_{2}n$. Higher polynomial degrees reduce approximation error but increase computational cost and multiplicative depth.
Limitation for General Computation: Polynomial approximation introduces inherent errors, especially near zero, making it unsuitable for applications requiring exact computation, such as genomics and finance. While it provides a form of computational generality, it is approximate rather than exact, limiting its applicability in error-sensitive domains. More details are in the Appendix A.1.
BGB/BFV with Polynomial Interpolation. In contrast to CKKS, computations in BGV/BFV are exact. Non-linear functions can be implemented via polynomial interpolation over the base field $\mathbb{F}_{p}$. For example, comparison operations can be expressed using Lagrange interpolation polynomials evaluated homomorphically on encrypted differences (narumanchi2017performance).
Limitation for General Computation: While polynomial interpolation supports exact non-linear operations, the degree of the interpolation polynomial grows exponentially with input bit-width when performed over $\mathbb{Z}_{p^{r}}$. To maintain efficiency, operations are typically confined to small primes $p\leq 257$, making this approach impractical for large inputs. Additionally, the plaintext space for linear operations ($\mathbb{Z}_{p^{d}}$) differs from that used for non-linear operations $\mathbb{F}_{p}$, preventing seamless composition without encoding conversion (as in Encoding Switching (zhang2023hebridge)). More details are in the Appendix A.2.
Polynomial Interpolation with Special Encoding. To enhance the scalability of polynomial interpolation, special encoding methods (iliashenko2021faster; tan2020efficient) decompose large integers into vectors of base-$p$ digits. Non-linear operations (e.g., comparisons) are then performed digit-wise, reducing depth from $\log_{2}{p^{r}}$ to $\log_{2}\log_{p}2^{b}+\log_{2}(p-1)+4$.
Limitation for General Computation: While these methods significantly improve the efficiency of non-linear operations, the specialized ciphertext format (vector encoding) does not support standard linear operations. Thus, they cannot seamlessly compose linear and non-linear operations without format conversion, disqualifying them as fully general FHE methods. More details are in the Appendix A.3.
Exponential Encoding (XCMP). XCMP (lu2018non) uses exponential encoding to perform private comparisons with constant multiplicative depth $O(1)$ by encoding values as polynomial degrees. This enables highly efficient comparisons but is limited to small input domains (typically ¡ $16$ bits) due to the polynomial degree constraint.
Limitation for General Computation: XCMP’s specialized encoding precludes direct composition with linear operations, and extensions to larger domains incur prohibitive multiplicative depth. More details are in the Appendix A.4.
General Functional Bootstrapping. Functional bootstrapping (alexandru2024general; lee2024functional; liu2023amortized) extends traditional bootstrapping by enabling the evaluation of arbitrary functions (via look-up tables) during noise refresh. Methods such as BFV-style (lee2024functional) and CKKS-style (alexandru2024general) functional bootstrapping approximate target functions using interpolation polynomials with complexity roughly proportional to $\sqrt{p}$ for first-order interpolation.
Limitation for General Computation: While theoretically powerful, functional bootstrapping remains practical only for small input spaces ($\leq 12$ bits) due to polynomial degree growth and scaling factor constraints. Moreover, extending to larger inputs requires digit-wise decomposition (similar to special encoding methods), which again precludes seamless linear-nonlinear composition. More details are in the Appendix A.5.
4. Relationship to Prior SoK work
Gouert et al. (gouert2023sok) presented a comprehensive SoK on FHE libraries, providing standardized benchmarks across major implementations including HElib, SEAL, Lattigo, OpenFHE, TFHE, and Concrete. Their work focuses on library-level performance evaluation—measuring runtime, memory consumption, ciphertext expansion, and parameter selection workflows across standard primitive operations (encryption, addition, multiplication, rotation). This library-focused SoK enables practitioners to compare implementation quality, identify performance bottlenecks, and select appropriate libraries for deployment. Additionally, Viand et al. (viand2021sok) primarily surveys FHE compilers, emphasizing performance, toolchain maturity, and developer accessibility. In contrast, our SoK addresses a fundamentally different research question: ”Can existing FHE methods support general AI computation?” Rather than benchmarking library implementations, we analyze the functional generality and computational scalability of underlying FHE methods when applied to AI workloads requiring both linear and non-linear operations. Table 2 summarizes the key distinctions between these three complementary systematization efforts.
Table 2. Comparison of FHE SoK contributions.
SoK Focus Units Evaluated Contribution
Viand et al. Compiler usability Compilers (EVA, CHET, Compiler landscape
(viand2021sok) & performance Cingulata, etc.) systematization
Gouert et al. Library Libraries (HElib, SEAL, Standardized library
(gouert2023sok) performance Lattigo, TFHE, PALISADE) benchmarking
Our Computational FHE Methods (TFHE, Scheme Method selection
Work generality Switching, Encoding Switching) for general AI
5. FHE Libraries Supporting General Computation